# Optimized Single-Core PCF-Based SPR Biosensor for High-Performance Early-Stage Multi-Cancer Detection

### v1

@@ I Introduction
Cancer remains one of the most critical global public health challenges, currently ranking as the second leading cause of mortality worldwide, surpassing even heart disease in some regions [1]. According to the World Health Organization (WHO), cancer is responsible for approximately one in six deaths globally, with an estimated 19.3 million new cases and 10 million deaths recorded in 2020 alone [2, 3]. Characterized by the uncontrolled proliferation of abnormal cells, cancer disrupts essential biological functions and often leads to fatal outcomes [1]. This unchecked cellular growth can be triggered by a range of factors, including exposure to toxic substances, radiation, fungal agents, poor dietary habits, obesity, excessive alcohol consumption, physical inactivity, and nutritional deficiencies [1]. Among the various forms of cancer, skin cancer and blood cancer are among the most frequently diagnosed types, with skin cancer showing a rising global incidence [1]. Despite significant advancements in medicine, a definitive cure for cancer remains elusive [1]. Therefore, early and accurate detection is paramount, as it substantially increases the chances of successful treatment and patient survival [4, 5]. Traditional diagnostic methods such as biopsies, physical exams, endoscopy, and imaging techniques like MRI and X-rays are often expensive and rely on bulky, complex equipment [5]. Consequently, there is an urgent demand for rapid, cost-effective, and highly sensitive diagnostic technologies that can facilitate early cancer detection [5].
@@
Optical biosensors have emerged as a leading solution for early disease detection due to their high sensitivity, compactness, rapid response, and low-cost fabrication. These advantages make them ideal for real-time biomedical applications such as cancer diagnosis. Various optical techniques, including fluorescence imaging, Raman spectroscopy, photothermal methods, terahertz (THz) spectroscopy, microfluidic devices, and fiber-optic sensors, have been explored for this purpose [6, 7, 8, 9]. Among these, Photonic Crystal Fiber (PCF)-based sensors stand out for their superior light-guiding properties and design flexibility. Unlike conventional fibers, PCFs feature periodic air-hole structures that provide enhanced control over dispersion, birefringence, and mode confinement. These properties allow PCFs to be tailored for ultra-sensitive detection of biological analytes. As a result, PCFs are widely used in sensing applications such as temperature, pressure, glucose, pH, DNA, and cancer biomarker detection. Their microstructured architecture enables strong interaction between light and analyte, leading to improved sensitivity and precision. With their compact design, low optical loss, and ability to perform label-free detection, PCF-based biosensors offer a powerful platform for next-generation, non-invasive cancer diagnostics.
@@
The development of biosensors, particularly surface plasmon resonance (SPR)-based platforms, has significantly advanced over recent decades. The earliest SPR biosensor, designed by Clark in 1985, was utilized for glucose detection in blood [10]. Since then, substantial progress has been made in applying SPR and photonic crystal fiber (PCF)-based sensors for biomedical diagnostics, particularly in cancer detection. Ahmet Yasli investigated SPR-based PCF biosensors [11], while other studies proposed optimized single-core [12] and dual-core PCF designs [13] for detecting various cancerous cells, including those affecting the breast, basal, and cervical tissues. A triple-core PCF biosensor was introduced for differentiating between normal and malignant blood cells based on refractive index (RI) contrast, offering early cancer detection capability[14]. Bulbul et al. explored PCF sensors in the terahertz (THz) regime using the finite element method for breast cancer detection [15]. Meanwhile, Rakhshani et al. developed a plasmonic sensor using a 3D nanorod metasurface, proposing a wide-angle absorber for cancer cell identification [16]. Recent innovations have focused on enhancing sensitivity and selectivity through novel material integration and hybrid configurations. For instance, Hemanth et al. introduced a 2D photonic crystal-based cervical cancer sensor achieving a peak sensitivity of 143 $nm/RIU$ [17]. Karki et al. made significant contributions to SPR sensor design by incorporating various layered materials such as black phosphorus (BP) [18], platinum diselenide $(PtSe_{2})$ and graphene [19], zinc sulfide (ZnS), silicon dioxide (SiO2) [20], and MXene [21] to enhance performance. Their sensors demonstrated high sensitivity (up to $374.31^{\circ}/\text{RIU}$) and figure of merit (FOM), though challenges in fabrication complexity and material stability were noted. Further, Karki and collaborators reviewed state-of-the-art SPR biosensing technologies for cancer detection and discussed avenues for improving sensitivity, specificity, and real-world applicability [22]. Kumar et al. developed a microfiber interferometer biosensor for ultra-low concentration detection [23], albeit without thoroughly addressing integration with current diagnostic frameworks. Pal et al. [24] and others validated long-range SPR sensors for blood-based biomarker detection. More recently, simulation-based designs have achieved remarkable sensitivities. Jabin et al. developed a D-shaped breast cancer sensor with peak sensitivity reaching 18,000 $nm/RIU$, surpassing most prior designs focused predominantly on basal cells [12].
@@
All fiber-based surface plasmon resonance (SPR) biosensors rely on the phase-matched coupling between the core-guided mode and the surface plasmon polariton (SPP) mode at the metal-dielectric interface. This coupling facilitates precise detection of refractive index (RI) variations in cellular samples by monitoring shifts in the resonance wavelength. When resonance conditions are satisfied, a distinct Gaussian-shaped peak emerges in the confinement loss spectrum, corresponding to the analyte’s specific RI[25]. The excitation of surface plasmons is highly dependent on the properties of the metal layer, which serves as the plasmonic medium. Common plasmonic materials include silver (Ag), copper (Cu), and gold (Au), among which gold is preferred due to its superior chemical stability, resistance to oxidation, biocompatibility, and ability to produce sharp and well-defined resonance peaks [26]. Photonic crystal fibers (PCFs) differ fundamentally from conventional optical fibers by incorporating a periodic array of air holes in the cladding, which induces spatial modulation of the refractive index and enhances light confinement within the core. These air holes are typically arranged in circular, hexagonal, or hybrid geometries, and can be engineered to tailor the modal properties of the fiber. In SPR-PCF biosensor designs, a section of the fiber is selectively infiltrated with the analyte, while a nanoscale metal film is strategically deposited at the core-cladding boundary or within inner air holes to facilitate plasmon excitation [27].
@@ II Design Strategy and Theoretical Modeling
The fabrication of this structure can be achieved using the stack-and-draw technique [28], a well-established method for PCF production. As shown in Fig. 1b, the preform is assembled by stacking thin-walled glass capillaries and solid rods in a pattern that replicates the desired fiber geometry. The U-shaped channel is then created using the polishing method described in [29]. Plasmonic materials, including gold and $V_{2}O_{5}$, are subsequently deposited onto the polished surface using chemical deposition techniques [30].
@@
Cladding radius
Air hole radius
Pitch
Gold layer
thickness
@@
$V_{2}O_{5}$
adhesive thickness
Perfectly matched
layer
@@
Silica is used as the background material of the proposed biosensor, and its refractive index (RI) is determined using the well-known Sellmeier equation, given by [31]:
@@
where $\lambda$ represents the incident light wavelength in $\mu m$, and $B_{n}$ and $C_{n}$ are the Sellmeier coefficients for n = 1, 2, 3. The values of $B_{1}$, $B_{2}$, and $B_{3}$ are 0.6961663, 0.4079426, and 0.8974794, respectively, while $C_{1}$, $C_{2}$, and $C_{3}$ are 0.068404, 0.1162414, and 9.896161 $\mu m$, as given in [32]. The frequency-dependent refractive index of the adhesive material, $V_{2}O_{5}$, is calculated using the following equation: [33]:
@@
where $\epsilon_{\infty}=5.9673$ is the permittivity at high frequency, $\Delta\epsilon=1.09$ is weighting factor, and $\omega$ is the angular frequency. The plasma frequency $\omega_{D}$ and damping frequency $\gamma_{D}$ are expressed by $\omega_{D}/{2\pi}=2113.6$ THz and $\gamma_{D}/{2\pi}=15.92$ THz, respectively. Frequency of the Lorentz oscillators and spectral width are given by $\Omega_{L}/{2\pi}=650.07$ THz and $\Gamma_{L}/{2\pi}=104.86$ THz, respectively [34].
@@ III-A Confinement Loss
where $k_{0}=\frac{2\pi}{\lambda}$ is the number of waves in free space, $\lambda(\mu m)$ is the operating wavelength and $Im(n_{eff})$ is the imaginary part of the effective refractive index. Fig. 3 presents the confinement loss (CL) spectra of the proposed PCF-SPR biosensor for differentiating between normal and cancerous cells in skin, blood, and adrenal gland tissues. In Fig. 3(a), for skin cells, the normal sample (n = 1.36) exhibits a minimum CL of approximately 3.21 $dB/cm$, while the cancerous counterpart (n = 1.38) shows a significantly higher peak CL of 17.67 $dB/cm$ near 2.2 $\mu m$, indicating strong plasmonic interaction. In Fig. 3(b), for blood cells, the normal cell (n = 1.376) demonstrates a low CL of about 13.2 $dB/cm$, whereas the cancerous cell (n = 1.39) reaches a peak CL of 138 $dB/cm$ around 2.8 $\mu m$, reflecting enhanced sensitivity. Similarly, in Fig. 3(c), for adrenal gland cells, the normal condition (n = 1.381) shows a CL minimum of approximately 7.9 $dB/cm$, while the cancerous cell (n = 1.395) exhibits a pronounced peak CL of 115.3 $dB/cm$ near 3.6 $\mu m$. These significant differences in peak confinement losses between normal and cancerous cells across all tissue types confirm the high sensitivity and diagnostic potential of the proposed PCF-SPR biosensor for effective cancer cell detection.
@@ III-D Figure of Merit (FOM)
where FWHM (in nm) represents the spectral width of the resonance curve at half its maximum loss value. A higher FOM indicates a sharper resonance and improved sensing performance, allowing for more accurate differentiation between analytes with similar refractive indices. Therefore, optimizing FOM is essential for enhancing the resolution, reliability, and diagnostic capability
Type of
Cancer
Type of
Cell
RI of
Normal Cell
RI of Cancer
Affected Cell
Peak loss
$(dB/cm)$
WS
$(nm/RIU)$
R
$(RIU)$
Skin Basal 1.360 1.380 17.67 15,000 $3.13\times 10^{-6}$
Blood Jurkat 1.376 1.390 138 53,571 $1.87\times 10^{-6}$
Adrenal PC12 1.381 1.395 115.3 98,214 $1.92\times 10^{-6}$
IV Structural Optimization
IV-A Airhole Optimization
Air holes in photonic crystal fiber (PCF) biosensors play a crucial role in controlling the propagation characteristics of the core-guided mode and facilitating efficient coupling with the surface plasmon polariton (SPP) mode at the metal-dielectric interface. The size and arrangement of these air holes directly influence the effective refractive index, confinement of the optical field, and overall sensitivity of the sensor. By optimizing the air hole radius (r), it is possible to finely tune the modal distribution and enhance phase-matching conditions, thereby improving plasmonic resonance and sensing performance. This section investigates the impact of varying air hole dimensions on the sensor’s performance to identify the optimal configuration that ensures maximum structural efficiency and sensing accuracy.
IV-A1 For skin cell
@@ IV-A3 For adrenal gland cell
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
@@
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
@@
IV-B Pitch $(\Lambda)$Optimization
The simulation was performed for two pitch values, $(\Lambda)$ = 6.0 $\mu m$ and $(\Lambda)$ = 6.2 $\mu m$, to evaluate their effect on the confinement loss and amplitude sensitivity of the proposed SPR-PCF biosensor. For skin cells, the confinement loss characteristics shown in Fig. 7(a) and Fig. 7(b) indicate that an increase in pitch $(\Lambda)$ from 6.0 $\mu m$ to 6.2 $\mu m$ results in higher peak confinement loss, while a smaller pitch enhances the amplitude sensitivity. This trend reflects a trade-off between loss and sensitivity, with $\Lambda$ = 6.0 $\mu m$ providing an optimal balance. A similar behavior is observed in the case of blood cells, as depicted in Fig. 7(c) and Fig. 7(d). Increasing the pitch leads to elevated confinement losses, while reducing the pitch improves amplitude sensitivity. The data in Table V and Table VI confirm that $\Lambda$ = 6.0 $\mu m$ offers the most favorable performance metrics for blood cell detection as well. For adrenal gland cells, Fig. 7(e) and Fig. 7(f) further reinforce this trend. As the pitch increases, the confinement loss rises, whereas a reduced pitch enhances the amplitude sensitivity. Again, the optimal pitch is found to be 6.0 $\mu m$, striking a practical compromise between maximizing sensitivity and minimizing loss. This establishes $\Lambda$ = 6.0 $\mu m$ as a robust design parameter for multi-type cancer cell detection using the proposed SPR-PCF biosensor.
@@ IV-B Pitch $(\Lambda)$Optimization
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
@@
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
@@
IV-C Optimizing Au Layer Thickness $(t_{Au})$
The effect of gold (Au) layer thickness on the performance of the SPR-PCF biosensor was analyzed by simulating two thicknesses: 40 $nm$ and 42 $nm$. Key performance metrics including wavelength sensitivity (WS), figure of merit (FOM), and resolution were evaluated for skin, blood, and adrenal gland cells, with results shown in Fig. 8(a)–Fig. 8(f) and Table VII and Table VIII. A gold thickness of 40 $nm$ consistently demonstrated superior performance for both skin and adrenal gland cells, yielding higher WS, improved FOM, and better resolution.
While the performance difference for blood cells was less pronounced, the 40 $nm$ layer still offered slightly better sensitivity and resolution, making it the preferred choice.
@@ IV-C Optimizing Au Layer Thickness $(t_{Au})$
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
@@
Type of
Cancer
Type of
cell
RI
Peak loss
$(dB/cm)$
SA
($RIU^{-1})$
SR
$(RIU)$
FOM
$(RIU^{-1})$
@@
Ref.
Year
Structure
RI range
Max. WS
$(nm/RIU)$
R
$(RIU)$
[35] 2023 Spiral Shaped 1.36-1.401 $2.33\times 10^{-4}$
[36] 2023 V-shaped 6,214 $1.66\times 10^{-5}$
[37] 2024 PCF based 1.36-1.392 6,250
[38] 2024 D-shaped 1.392-1.399 12,857

### v3

@@ I Introduction
Cancer remains one of the most critical global public health challenges, currently ranking as the second leading cause of mortality worldwide, surpassing even heart disease in some regions [siegel2017colorectal]. According to the World Health Organization (WHO), cancer is responsible for approximately one in six deaths globally, with an estimated 19.3 million new cases and 10 million deaths recorded in 2020 alone [singh2019nanotechnology, sung2021global]. Characterized by the uncontrolled proliferation of abnormal cells, cancer disrupts essential biological functions and often leads to fatal outcomes [siegel2017colorectal]. This unchecked cellular growth can be triggered by a range of factors, including exposure to toxic substances, radiation, fungal agents, poor dietary habits, obesity, excessive alcohol consumption, physical inactivity, and nutritional deficiencies [siegel2017colorectal]. Among the various forms of cancer, skin cancer and blood cancer are among the most frequently diagnosed types, with skin cancer showing a rising global incidence [siegel2017colorectal]. Despite significant advancements in medicine, a definitive cure for cancer remains elusive [siegel2017colorectal]. Therefore, early and accurate detection is paramount, as it substantially increases the chances of successful treatment and patient survival [liang2007determining, sharma2015photonic]. Traditional diagnostic methods such as biopsies, physical exams, endoscopy, and imaging techniques like MRI and X-rays are often expensive and rely on bulky, complex equipment [sharma2015photonic]. Consequently, there is an urgent demand for rapid, cost-effective, and highly sensitive diagnostic technologies that can facilitate early cancer detection [sharma2015photonic].
@@
Optical biosensors have emerged as a leading solution for early disease detection due to their high sensitivity, compactness, rapid response, and low-cost fabrication. These advantages make them ideal for real-time biomedical applications such as cancer diagnosis. Various optical techniques, including fluorescence imaging, Raman spectroscopy, photothermal methods, terahertz (THz) spectroscopy, microfluidic devices, and fiber-optic sensors, have been explored for this purpose [zaytsev2015vivo, hajba2014circulating, liu2014surface, li2013detection]. Among these, Photonic Crystal Fiber (PCF)-based sensors stand out for their superior light-guiding properties and design flexibility. Unlike conventional fibers, PCFs feature periodic air-hole structures that provide enhanced control over dispersion, birefringence, and mode confinement. These properties allow PCFs to be tailored for ultra-sensitive detection of biological analytes. As a result, PCFs are widely used in sensing applications such as temperature, pressure, glucose, pH, DNA, and cancer biomarker detection. Their microstructured architecture enables strong interaction between light and analyte, leading to improved sensitivity and precision. With their compact design, low optical loss, and ability to perform label-free detection, PCF-based biosensors offer a powerful platform for next-generation, non-invasive cancer diagnostics.
@@
The development of biosensors, particularly surface plasmon resonance (SPR)-based platforms, has significantly advanced over recent decades. The earliest SPR biosensor, designed by Clark in 1985, was utilized for glucose detection in blood [hasan2023indium]. Since then, substantial progress has been made in applying SPR and photonic crystal fiber (PCF)-based sensors for biomedical diagnostics, particularly in cancer detection. Ahmet Yasli investigated SPR-based PCF biosensors [yasli2021cancer], while other studies proposed optimized single-core [jabin2019surface] and dual-core PCF designs [ayyanar2018photonic] for detecting various cancerous cells, including those affecting the breast, basal, and cervical tissues. A triple-core PCF biosensor was introduced for differentiating between normal and malignant blood cells based on refractive index (RI) contrast, offering early cancer detection capability[gong2024lab]. Bulbul et al. explored PCF sensors in the terahertz (THz) regime using the finite element method for breast cancer detection [bulbul2020design]. Meanwhile, Rakhshani et al. developed a plasmonic sensor using a 3D nanorod metasurface, proposing a wide-angle absorber for cancer cell identification [rakhshani2021wide]. Recent innovations have focused on enhancing sensitivity and selectivity through novel material integration and hybrid configurations. For instance, Hemanth et al. introduced a 2D photonic crystal-based cervical cancer sensor achieving a peak sensitivity of 143 $nm/RIU$ [kumar20202d]. Karki et al. made significant contributions to SPR sensor design by incorporating various layered materials such as black phosphorus (BP) [karki2025tuning], platinum diselenide $(PtSe_{2})$ and graphene [karki2024platinum], zinc sulfide (ZnS), silicon dioxide (SiO2) [karki2022zinc], and MXene [karki2024gold] to enhance performance. Their sensors demonstrated high sensitivity (up to $374.31^{\circ}/\text{RIU}$) and figure of merit (FOM), though challenges in fabrication complexity and material stability were noted. Further, Karki and collaborators reviewed state-of-the-art SPR biosensing technologies for cancer detection and discussed avenues for improving sensitivity, specificity, and real-world applicability [karki2022advances]. Kumar et al. developed a microfiber interferometer biosensor for ultra-low concentration detection [kumar2019ultrasensitive], albeit without thoroughly addressing integration with current diagnostic frameworks. Pal et al. [pal2024highly] and others validated long-range SPR sensors for blood-based biomarker detection. More recently, simulation-based designs have achieved remarkable sensitivities. Jabin et al. developed a D-shaped breast cancer sensor with peak sensitivity reaching 18,000 $nm/RIU$, surpassing most prior designs focused predominantly on basal cells [jabin2019surface].
@@
All fiber-based surface plasmon resonance (SPR) biosensors rely on the phase-matched coupling between the core-guided mode and the surface plasmon polariton (SPP) mode at the metal-dielectric interface. This coupling facilitates precise detection of refractive index (RI) variations in cellular samples by monitoring shifts in the resonance wavelength. When resonance conditions are satisfied, a distinct Gaussian-shaped peak emerges in the confinement loss spectrum, corresponding to the analyte’s specific RI[malakar2025performance]. The excitation of surface plasmons is highly dependent on the properties of the metal layer, which serves as the plasmonic medium. Common plasmonic materials include silver (Ag), copper (Cu), and gold (Au), among which gold is preferred due to its superior chemical stability, resistance to oxidation, biocompatibility, and ability to produce sharp and well-defined resonance peaks [sultana2021highly]. Photonic crystal fibers (PCFs) differ fundamentally from conventional optical fibers by incorporating a periodic array of air holes in the cladding, which induces spatial modulation of the refractive index and enhances light confinement within the core. These air holes are typically arranged in circular, hexagonal, or hybrid geometries, and can be engineered to tailor the modal properties of the fiber. In SPR-PCF biosensor designs, a section of the fiber is selectively infiltrated with the analyte, while a nanoscale metal film is strategically deposited at the core-cladding boundary or within inner air holes to facilitate plasmon excitation [haque2021highly].
@@ II Design Strategy and Theoretical Modeling
The fabrication of this structure can be achieved using the stack-and-draw technique [rifat2017highly], a well-established method for PCF production. As shown in Fig. 1b, the preform is assembled by stacking thin-walled glass capillaries and solid rods in a pattern that replicates the desired fiber geometry. The U-shaped channel is then created using the polishing method described in [wu2018ultrahigh]. Plasmonic materials, including gold and $V_{2}O_{5}$, are subsequently deposited onto the polished surface using chemical deposition techniques [chen2018surface].
@@
Cladding radius Air hole radius Pitch Gold layer thickness
@@
$V_{2}O_{5}$ adhesive thickness Perfectly matched layer
@@
Silica is used as the background material of the proposed biosensor, and its refractive index (RI) is determined using the well-known Sellmeier equation, given by [hassan2020liquid5]:
@@
where $\lambda$ represents the incident light wavelength in $\mu m$, and $B_{n}$ and $C_{n}$ are the Sellmeier coefficients for n = 1, 2, 3. The values of $B_{1}$, $B_{2}$, and $B_{3}$ are 0.6961663, 0.4079426, and 0.8974794, respectively, while $C_{1}$, $C_{2}$, and $C_{3}$ are 0.068404, 0.1162414, and 9.896161 $\mu m$, as given in [ullah2024highly6]. The frequency-dependent refractive index of the adhesive material, $V_{2}O_{5}$, is calculated using the following equation: [atuchin2014effects]:
@@
where $\epsilon_{\infty}=5.9673$ is the permittivity at high frequency, $\Delta\epsilon=1.09$ is weighting factor, and $\omega$ is the angular frequency. The plasma frequency $\omega_{D}$ and damping frequency $\gamma_{D}$ are expressed by $\omega_{D}/{2\pi}=2113.6$ THz and $\gamma_{D}/{2\pi}=15.92$ THz, respectively. Frequency of the Lorentz oscillators and spectral width are given by $\Omega_{L}/{2\pi}=650.07$ THz and $\Gamma_{L}/{2\pi}=104.86$ THz, respectively [zhu2017surface8].
@@ III-A Confinement Loss
where $k_{0}=\frac{2\pi}{\lambda}$ is the number of waves in free space, $\lambda(\mu m)$ is the operating wavelength and $Im(n_{eff})$ is the imaginary part of the effective refractive index.
Fig. 3 presents the confinement loss (CL) spectra of the proposed PCF-SPR biosensor for differentiating between normal and cancerous cells in skin, blood, and adrenal gland tissues. In Fig. 3(a), for skin cells, the normal sample (n = 1.36) exhibits a minimum CL of approximately 3.21 $dB/cm$, while the cancerous counterpart (n = 1.38) shows a significantly higher peak CL of 17.67 $dB/cm$ near 2.2 $\mu m$, indicating strong plasmonic interaction. In Fig. 3(b), for blood cells, the normal cell (n = 1.376) demonstrates a low CL of about 13.2 $dB/cm$, whereas the cancerous cell (n = 1.39) reaches a peak CL of 138 $dB/cm$ around 2.8 $\mu m$, reflecting enhanced sensitivity. Similarly, in Fig. 3(c), for adrenal gland cells, the normal condition (n = 1.381) shows a CL minimum of approximately 7.9 $dB/cm$, while the cancerous cell (n = 1.395) exhibits a pronounced peak CL of 115.3 $dB/cm$ near 3.6 $\mu m$. These significant differences in peak confinement losses between normal and cancerous cells across all tissue types confirm the high sensitivity and diagnostic potential of the proposed PCF-SPR biosensor for effective cancer cell detection.
@@
Type of Cancer Type of Cell RI of Normal Cell RI of Cancer Affected Cell Peak loss $(dB/cm)$ WS $(nm/RIU)$ R $(RIU)$
Skin Basal 1.360 1.380 17.67 15,000 $3.13\times 10^{-6}$
Blood Jurkat 1.376 1.390 138 53,571 $1.87\times 10^{-6}$
Adrenal PC12 1.381 1.395 115.3 98,214 $1.92\times 10^{-6}$
@@ III-D Figure of Merit (FOM)
where FWHM (in nm) represents the spectral width of the resonance curve at half its maximum loss value. A higher FOM indicates a sharper resonance and improved sensing performance, allowing for more accurate differentiation between analytes with similar refractive indices. Therefore, optimizing FOM is essential for enhancing the resolution, reliability, and diagnostic capability.
@@
IV Structural Optimization
@@
IV-A Airhole Optimization
Air holes in photonic crystal fiber (PCF) biosensors play a crucial role in controlling the propagation characteristics of the core-guided mode and facilitating efficient coupling with the surface plasmon polariton (SPP) mode at the metal-dielectric interface. The size and arrangement of these air holes directly influence the effective refractive index, confinement of the optical field, and overall sensitivity of the sensor.
By optimizing the air hole radius (r), it is possible to finely tune the modal distribution and enhance phase-matching conditions, thereby improving plasmonic resonance and sensing performance. This section investigates the impact of varying air hole dimensions on the sensor’s performance to identify the optimal configuration that ensures maximum structural efficiency and sensing accuracy.
IV-A1 For skin cell
IV-A2 For blood cell
IV-A3 For adrenal gland cell
@@
Type of Cancer Type of cell RI Peak loss $(dB/cm)$ SA ($RIU^{-1})$ SR $(RIU)$ FOM $(RIU^{-1})$
@@
Ref. Year Structure RI range Max. WS $(nm/RIU)$ R $(RIU)$
[mittal2023spiral] 2023 Spiral Shaped 1.36-1.401 $2.33\times 10^{-4}$
[abdelghaffar2023highly] 2023 V-shaped 6,214 $1.66\times 10^{-5}$
[jain2024pcf] 2024 PCF based 1.36-1.392 6,250
[nagavel2024highly] 2024 D-shaped 1.392-1.399 12,857
@@
IV-B Pitch $(\Lambda)$Optimization
The simulation was performed for two pitch values, $(\Lambda)$ = 6.0 $\mu m$ and $(\Lambda)$ = 6.2 $\mu m$, to evaluate their effect on the confinement loss and amplitude sensitivity of the proposed SPR-PCF biosensor. For skin cells, the confinement loss characteristics shown in Fig. 7(a) and Fig. 7(b) indicate that an increase in pitch $(\Lambda)$ from 6.0 $\mu m$ to 6.2 $\mu m$ results in higher peak confinement loss, while a smaller pitch enhances the amplitude sensitivity. This trend reflects a trade-off between loss and sensitivity, with $\Lambda$ = 6.0 $\mu m$ providing an optimal balance. A similar behavior is observed in the case of blood cells, as depicted in Fig. 7(c) and Fig. 7(d). Increasing the pitch leads to elevated confinement losses, while reducing the pitch improves amplitude sensitivity. The data in Table V and Table VI confirm that $\Lambda$ = 6.0 $\mu m$ offers the most favorable performance metrics for blood cell detection as well. For adrenal gland cells, Fig. 7(e) and Fig. 7(f) further reinforce this trend. As the pitch increases, the confinement loss rises, whereas a reduced pitch enhances the amplitude sensitivity. Again, the optimal pitch is found to be 6.0 $\mu m$, striking a practical compromise between maximizing sensitivity and minimizing loss. This establishes $\Lambda$ = 6.0 $\mu m$ as a robust design parameter for multi-type cancer cell detection using the proposed SPR-PCF biosensor, making it a preferred choice
IV-C Optimizing Au Layer Thickness $(t_{Au})$
The effect of gold (Au) layer thickness on the performance of the SPR-PCF biosensor was analyzed by simulating two thicknesses: 40 $nm$ and 42 $nm$. Key performance metrics including wavelength sensitivity (WS), figure of merit (FOM), and resolution were evaluated for skin, blood, and adrenal gland cells, with results shown in Fig. 8(a)–Fig. 8(f) and Table VII and Table VIII. A gold thickness of 40 $nm$ consistently demonstrated superior performance for both skin and adrenal gland cells, yielding higher WS, improved FOM, and better resolution. While the performance difference for blood cells was less pronounced, the 40 $nm$ layer still offered slightly better sensitivity and resolution.